### v2

@@ A.3 Proof of Proposition 1
We know from Corollary 1 that all voters with $s>0$ have belief $\pi(s)=\hat{\pi}$, while those with $s<0$ have belief $\pi(s)=0$. Hence, the latter always vote for policy 0. Wether a voter with $s>0$ votes for policy 1 or policy 0 depends on the comparison of $\tilde{\pi}$ and $\hat{\pi}$. In particular, we know from (5) that a voter votes for policy 1 if $\pi(s)>\tilde{\pi}$, for policy 0 if $\pi(s)<\tilde{\pi}$, and they choose each with a probability of 50% if $\pi(s)=\tilde{\pi}$. Hence, if $\hat{\pi}<\tilde{\pi}$, all voters vote for policy 0. If $\hat{\pi}=\tilde{\pi}$, all voters receiving $s<0$ plus half of the voter receiving $s>0$ vote for policy 0, while the other half of $s>0$ voters vote for policy 1. As a consequence, policy 0 wins all the time if $\hat{\pi}\leq\tilde{\pi}$. This implies that we must have $\kappa^{*}(s,\mathbf{p},\tilde{\mu}^{*})=0$ for all $s$.
@@
However, if $\hat{\pi}>\tilde{\pi}$, all voters vote informatively, and thus the optimal policy always wins. It follows that in state 0 a majority votes for policy 0, while in state 1 a majority votes for policy 1. Hence, voters know that the policy that matches the state wins all the time. The belief about the state is $\pi(s)=\hat{\pi}$ if $s\geq 0$, implying $\left.\kappa^{*}(s,\mathbf{p},\tilde{\mu}^{*})\right|_{s\geq 0}=\hat{\pi}$, and $\pi(s)=0$ if $s<0$, and hence $\left.\kappa^{*}(s,\mathbf{p},\tilde{\mu}^{*})\right|_{s<0}=0$.
Could there be other pure strategy equilibria? $\tilde{\mu}^{*}$ is uniquely determined, and hence also $\pi(s,\mathbf{p},\tilde{\mu}^{*})$ is uniquely determined. But this implies that equilibrium vote shares are unique, leaving no room for beliefs $\kappa(s,\mathbf{p},\tilde{\mu})$ that are consistent with these vote shares and that differ from the ones established before. Hence, if $\Delta>1$, no other equilibrium in the voting subgame can exist.
$\Delta\leq 1$:
@@
First assume that $\kappa(s,\mathbf{p},\tilde{\mu})$ is increasing in $s$ in the sense that $\text{Sign}\left[\kappa(s,\mathbf{p},\tilde{\mu})-\tilde{\kappa}\right]=\text{Sign}\left[s\right]$. Then any voter with $s<0$ chooses $\tilde{\mu}^{*}=\infty$, and thus all these voters hold belief $\pi(s,\mathbf{p},\tilde{\mu}^{*})=0$ and vote for policy 0. Moreover, any voter with $s>0$ chooses also $\tilde{\mu}^{*}=\infty$, and thus all these voters hold belief $\pi(s,\mathbf{p},\tilde{\mu}^{*})=1$ and vote for policy 1. This implies that indeed the policy that matches the state always wins and we need to have $\left.\kappa(s,\mathbf{p},\tilde{\mu}^{*})\right|_{s<0}=0$ as well as $\left.\kappa(s,\mathbf{p},\tilde{\mu}^{*})\right|_{s>0}=1$, proving part 2 (a) of the proposition.
@@
Next, assume voters generally believe that policy 0 will be implemented, independent of the true state $\omega$. Then $\kappa(s,\mathbf{p},\tilde{\mu}^{*})=0$ for all $s$. Then any voter with $s<0$ chooses $\tilde{\mu}^{*}=\infty$, and thus all these voters hold belief $\pi(s,\mathbf{p},\tilde{\mu}^{*})=0$ and vote for policy 0. Moreover, any voter with $s>0$ chooses $\tilde{\mu}^{*}=0$. If and only if $\hat{\pi}\leq\tilde{\pi}$, this implies that indeed always a majority of voters votes for policy 0, and hence when $\hat{\pi}\leq\tilde{\pi}$ there exists an equilibrium of the voting subgame in which policy 0 is always chosen. This proves part 2 (b).
@@
Now assume voters generally believe that policy 1 will be implemented, independent of the true state $\omega$. Then $\kappa(s,\mathbf{p},\tilde{\mu}^{*})=1$ for all $s$, and any voter with $s>0$ chooses $\tilde{\mu}^{*}=\infty$, and thus all these voters hold belief $\pi(s,\mathbf{p},\tilde{\mu}^{*})=1$ and vote for policy 1. Moreover, any voter with $s<0$ chooses $\tilde{\mu}^{*}=0$. If and only if $\hat{\pi}\geq\tilde{\pi}$, this implies that indeed always a majority of voters votes for policy 1, and hence when $\hat{\pi}\geq\tilde{\pi}$ there exists an equilibrium of the voting subgame in which policy 1 is always chosen. This proves part 2 (c).
Consider a situation in which beliefs are just as in the equilibrium discussed above for the case of $\Delta>1$. That is, $\kappa(s,\mathbf{p},\tilde{\mu})=0$ for all $s$. Then, as before, $\tilde{\mu}^{*}(s)=\infty$ and $\pi(s,\mathbf{p},\tilde{\mu}^{*})=0$ if $s<0$, and $\tilde{\mu}^{*}(s)=0$ and $\pi(s,\mathbf{p},\tilde{\mu}^{*})=q$ else. It follows that if $q\leq\frac{1}{1+\beta}$, a majority of voters always votes for policy 0. Hence, this is indeed an equilibrium.
Consider next a situation in which beliefs are just opposite to what we have seen above, namely $\kappa(s,\mathbf{p},\tilde{\mu})=1$ for all $s$. Then, $\tilde{\mu}^{*}(s)=0$ and $\pi(s,\mathbf{p},\tilde{\mu}^{*})=1$ if $s>0$, and $\tilde{\mu}^{*}(s)=0$ and $\pi(s,\mathbf{p},\tilde{\mu}^{*})=q$ else. It follows that if $q\geq\frac{1}{1+\beta}$, a majority of voters always votes for policy 1. Hence, this is also an equilibrium.
@@
Finally, consider part 2 (d). An equilibrium in which the optimal policy never wins would imply policy 1 wins if $\omega=0$ and policy 0 wins if $\omega=1$. Hence, we would need to have $\kappa(s,\mathbf{p},\tilde{\mu})$ weakly decreasing in $s$. Consider a voter with signal $s>0$. It must be true that most of these voters vote for policy 0. Hence, it cannot be true that $\kappa\geq\tilde{\kappa}$. If $\kappa(s,\mathbf{p},\tilde{\mu})<\tilde{\kappa}$, these voters all hold belief $\hat{\pi}$. Next consider voters with $s<0$. It must be true that most of these voters vote for policy 1, implying $\kappa(s,\mathbf{p},\tilde{\mu})\leq\tilde{\kappa}$ is not possible. Thus, because $\kappa(s,\mathbf{p},\tilde{\mu})>\tilde{\kappa}$, these voters all also hold belief $\hat{\pi}$. If $\hat{\pi}>\tilde{\pi}$, policy 1 always wins, contradicting that the optimal policy never wins. If $\hat{\pi}<\tilde{\pi}$, policy 0 always wins, also contradicting that the optimal policy never wins. Finally, if $\hat{\pi}=\tilde{\pi}$, in each state, each policy wins with 50%, and thus also the optimal policy wins with a chance of 50%. This implies that the actual state is irrelevant for the probability of each policy winning, and thus we must have $\kappa(s,\mathbf{p},\tilde{\mu})=\frac{1}{2}$ for all $s$. But then it is not possible that $\kappa(s,\mathbf{p},\tilde{\mu})$ differs for voters with $s>0$ and voters with $s<0$. Hence, such an equilibrium of the voting subgame cannot exist.
This complete the proof of the proposition.∎
@@ A.5 Proof of Proposition 3
We first consider the case of severe climate change being catastrophic, $\Delta>1$. The other case of severe climate change having only mild baseline welfare consequences will be considered thereafter. Note that $\hat{\pi}=q\leq\tilde{\pi}$ in a symmetric pure strategy equilibrium. Suppose that off the equilibrium path voters observe $\tilde{\mathbf{p}}\in\{\{0,1\},\{1,0\}\}$.
$\Delta>1$:
@@
We know from Corollary 1 that $\pi(s,\tilde{\mathbf{p}},\tilde{\mu}^{*})=0$ if $s<0$ and $\pi(s,\tilde{\mathbf{p}},\tilde{\mu}^{*})=\hat{\pi}$ if $s\geq 0$. Hence, $\hat{\pi}\leq\tilde{\pi}$, and thus no voter ever votes for policy 1, whereas policy 0 always wins at the ballot. It follows that $\kappa^{*}(s,\tilde{\mathbf{p}},\tilde{\mu}^{*})=0$, with $\tilde{\mu}^{*}(s)$ as described in Lemma 1. Thus, the unique equilibrium is candidates choosing $\mathbf{p}^{*}=(0,0)$. This is supported by $\kappa^{*}(s,\mathbf{p}^{*},\tilde{\mu}^{*})=0$ and $\tilde{\mu}^{*}(s)$ as described in Lemma 1.
$\Delta\leq 1$:
Assume first $\kappa(s,\tilde{\mathbf{p}},\tilde{\mu})=0$ for all $s$. Then it follows from Corollary 1 that $\pi(s,\tilde{\mathbf{p}},\tilde{\mu})=0$ if $s<0$ and $\pi(s,\tilde{\mathbf{p}},\tilde{\mu})=\hat{\pi}=q$ if $s>0$. Because $q\leq\frac{1}{1+\beta}$, a majority of voters always votes for policy 0. It follows from Proposition 2 that no candidate has an incentive to campaign on $p=1$. Consequently, $\mathbf{p}^{*}=(0,0)$, $\kappa(s,\mathbf{p}^{*},\tilde{\mu}^{*})=0$, and $\tilde{\mu}^{*}$ follows from Lemma 1.
Suppose next that $\Delta<1$ and that $\kappa>\tilde{\kappa}$ if $s>0$ and $\kappa<\tilde{\kappa}$ if $s<0$. Note that this means that we must have $\Delta<1$. Then it follows from Corollary 1 that $\pi(s,\tilde{\mathbf{p}},\tilde{\mu})=0$ if $s<0$ and $\pi(s,\tilde{\mathbf{p}},\tilde{\mu})=1$ if $s>0$. Hence, all voters all the time vote informatively, and thus the policy that matches the state wins, implying $\kappa=\pi(s,\tilde{\mathbf{p}},\tilde{\mu})$. It follows from Proposition 2 that, if candidates expect voters to behave this way, they have an incentive to choose the socially optimal policies. Hence, $\mathbf{p}^{*}=(\omega,\omega)$, $\kappa(s,\mathbf{p}^{*},\tilde{\mu}^{*})=\pi(s,\mathbf{p}^{*},\tilde{\mu}^{*})$, and $\tilde{\mu}^{*}$ follows from Lemma 1.
Now suppose that $\Delta=1$. Moreover, suppose an efficient equilibrium exists also in this case, with $\kappa(s,\mathbf{p},\tilde{\mu})=1=\tilde{\kappa}$ if $s>0$, and $\kappa=0<\tilde{\kappa}$ if $s<0$. Then it follows from Corollary 1 that $\pi(s,\tilde{\mathbf{p}},\tilde{\mu})=0$ if $s<0$, and thus any voter with $s<0$ votes for policy 0. For voters with $s>0$, $\tilde{\mu}^{*}=\mu$, and thus these voters hold perfect Bayesian beliefs. This implies, by Assumption 2, that voters vote informatively. However, then the right policy is chosen in any state, and thus we need to have $\kappa(s,\mathbf{p},\tilde{\mu})=\pi(s,\mathbf{p},\tilde{\mu})$. But $\pi(s,\mathbf{p},\tilde{\mu})<1$, and therefore voters with $s>0$ would choose $\tilde{\mu}=0$ instead of $\tilde{\mu}=\mu$. Hence, when $\Delta=1$, no efficient equilibrium exists.
@@
It follows from Proposition 1 part 2 (d) that no equilibrium can exist in which always the wrong policy is chosen. Hence, we now only need to show that no equilibrium, in which policy 1 always wins, can exist. Suppose to the contrary that voters believe that policy 1 wins with certainty, $\kappa(s,\mathbf{p},\tilde{\mu})=1\geq\tilde{\kappa}$ for all $s$. Then it follows from Corollary 1 that, if $\Delta<1$, $\pi(s,\tilde{\mathbf{p}},\tilde{\mu})=q\leq\tilde{\pi}$ if $s<0$ and $\pi(s,\tilde{\mathbf{p}},\tilde{\mu})=1$ if $s>0$. Therefore, all voters with $s>0$ always vote for policy 1. If sufficiently many voters with $s<0$ also vote for policy 1, policy 1 wins all the time. This is the case iff $q\geq\frac{1}{1+\beta}$. However, this contradicts Assumption 1. A similar argument shows that $\mathbf{p}=(1,1)$ cannot be equilibrium, either, if $\Delta=1$. Hence, it follows that no equilibrium, in which policy 1 is always chosen, exists. This proves the proposition. ∎
A.6 Proof of Lemma 2
Take the belief about $\omega$ of a voter who receives $s$ and chooses $\tilde{\alpha}$:
$\pi(s,\tilde{\alpha})=\frac{q\left(\tilde{\alpha}g(s)+(1-\tilde{\alpha})f_{1}(s)\right)}{q\left(\tilde{\alpha}g(s)+(1-\tilde{\alpha})f_{1}(s)\right)+(1-q)\left(\tilde{\alpha}g(s)+(1-\tilde{\alpha})f_{0}(s)\right)}.$
Anticipatory utility is defined as before in (2). Hence, it’s derivative with respect to $\tilde{\alpha}$ is
$\dfrac{\partial AU}{\partial\tilde{\alpha}}=\frac{(1-q)qg(s)\left[\beta+\Delta-(\beta+1)\kappa\right](f_{1}(s)-f_{0}(s))}{\left[q\left(\tilde{\alpha}g(s)+(1-\tilde{\alpha})f_{1}(s)\right)+(1-q)\left(\tilde{\alpha}g(s)+(1-\tilde{\alpha})f_{0}(s)\right)\right]^{2}}$
The derivative is positive if $f_{1}(s)-f_{0}(s)>0\Leftrightarrow s>0$, negative if $f_{1}(s)-f_{0}(s)<0\Leftrightarrow s<0$, and zero else. This proves the result. ∎
A.7 Proof of Proposition 4
It follows from Lemma 2 that independent of the true state $\omega$, a majority of voters prefers to vote for policy 0 when offered the choice. The proof then follows from Proposition 2. ∎
A.8 Proof of Lemma 3
We focus on pure strategy equilibria. It then follows from Lemma 1 that when $s>0$, a voter would either like to choose $\tilde{\mu}=\infty$, or $\tilde{\mu}=\mu$. Moreover, when $s<0$, a voter would either like to choose $\tilde{\mu}=0$, or $\tilde{\mu}=\mu$. Next have a look how a voter’s objective function changes with $s$:
$\frac{\partial W(s,\mathbf{p},\tilde{\mu})}{\partial s}=-\frac{2\mu(1-q)q\left[\beta+\Delta-(\beta+1)\kappa(s,\mathbf{p},\tilde{\mu})\right]e^{\frac{2\mu s}{\tilde{\mu}^{2}}}}{\left(q\tilde{\mu}\left(e^{\frac{2\mu s}{\tilde{\mu}^{2}}}-1\right)+\tilde{\mu}\right)^{2}}<0,$ (7)
and hence $W(s,\mathbf{p},\tilde{\mu})$ strictly decreases in $s$ for all $\tilde{\mu}$. Note that when $\tilde{\mu}=0$, then the realization of $s$ is irrelevant as the belief is $\left.\pi(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=0\wedge s>0}=q$. Hence, the voter with the greatest incentive to distort $\mu$ has $s\rightarrow\infty$. For this voter, the payoff of not distorting and instead choosing $\tilde{\mu}=\mu$ is
$\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=\mu\wedge s\rightarrow\infty}=-\beta(1-\kappa(s,\mathbf{p},\tilde{\mu}))-\Delta.$
If, to the contrary, any such voter chooses $\tilde{\mu}=0$, the payoff is
$\left.W(s,\mathbf{p},\tilde{\mu})\right|_{\tilde{\mu}=0\wedge s>0}=-\kappa(s,\mathbf{p},\tilde{\mu})(1-q)-q(\beta(1-\kappa(s,\mathbf{p},\tilde{\mu}))+\Delta)-\gamma.$
If $\gamma\geq\gamma^{+}\equiv(1-q)(\beta(1-\kappa(s,\mathbf{p},\tilde{\mu}))+\Delta-\kappa(s,\mathbf{p},\tilde{\mu}))$, then no voter with $s>0$ has an incentive to distort the issue’s complexity. Otherwise, that is, if $\gamma<\gamma^{+}$, then there exists a unique signal realization $s=s^{+}>0$ such that a voter receiving this signal is indifferent between choosing $\tilde{\mu}=0$ and $\tilde{\mu}=\mu$. The payoff of a voter choosing to not distort a signal $s$ is

### v5

@@ A.3 Proof of Proposition 1
We know from Corollary 1 that all voters with $s>0$ have belief $\pi(s,\mathbf{p},\tilde{\mu})=\hat{\pi}$, while those with $s<0$ have belief $\pi(s,\mathbf{p},\tilde{\mu})=0$. Hence, the latter always vote for policy 0. whether a voter with $s>0$ votes for policy 1 or policy 0 depends on the comparison of $\tilde{\pi}$ and $\hat{\pi}$. In particular, we know from (5) that a voter votes for policy 1 if $\pi(s,\mathbf{p},\tilde{\mu})>\tilde{\pi}$, for policy 0 if $\pi(s,\mathbf{p},\tilde{\mu})<\tilde{\pi}$, and they choose each with a probability of 50% if $\pi(s,\mathbf{p},\tilde{\mu})=\tilde{\pi}$. Hence, if $\hat{\pi}<\tilde{\pi}$, all voters vote for policy 0. If $\hat{\pi}=\tilde{\pi}$, all voters receiving $s<0$ plus half of the voter receiving $s>0$ vote for policy 0, while the other half of $s>0$ voters vote for policy 1. As a consequence, policy 0 wins all the time if $\hat{\pi}\leq\tilde{\pi}$. This implies that we must have $\kappa^{*}=0$ for all $s$.
@@
However, if $\hat{\pi}>\tilde{\pi}$, all voters vote informatively, and thus the optimal policy always wins. It follows that in state 0 a majority votes for policy 0, while in state 1 a majority votes for policy 1. Hence, voters know that the policy that matches the state wins all the time. The belief about the state is $\pi(s,\mathbf{p},\tilde{\mu})=\hat{\pi}$ if $s\geq 0$, implying $\left.\kappa^{*}\right|_{s\geq 0}=\hat{\pi}$, and $\pi(s,\mathbf{p},\tilde{\mu})=0$ if $s<0$, and hence $\left.\kappa^{*}\right|_{s<0}=0$.
[…]
$\Delta=1$:
If $\kappa=1$, then $\pi(s,\mathbf{p},\tilde{\mu})=\pi(s,\mathbf{p},\mu)$. Hence, any voter with $s<0$ will always vote for policy 0. This means policy 0 wins at least if $\omega=0$, contradicting $\kappa=1$. Hence, if $\Delta=1$, there cannot be an equilibrium with $\kappa=1$.
If $\kappa<1=\tilde{\kappa}$, then $\tilde{\mu}^{*}=0$ if $s>0$ and $\tilde{\mu}^{*}$$\infty$ if $s<0$. It follows that equilibrium is the same as in the case of $\Delta>1$.
$\Delta<1$:
@@
First assume that $\kappa$ positively correlates with $s$ in the sense that $\text{Sign}\left[\kappa-\tilde{\kappa}\right]=\text{Sign}\left[s\right]$. Any voter with $s<0$ chooses $\tilde{\mu}^{*}$$\infty$, and thus all these voters hold belief $\pi(s,\mathbf{p},\tilde{\mu}^{*})=0$ and vote for policy 0. Moreover, any voter with $s>0$ chooses also $\tilde{\mu}^{*}$$\infty$, and thus all these voters hold belief $\pi(s,\mathbf{p},\tilde{\mu}^{*})=1$ and vote for policy 1. This implies that indeed the policy that matches the state always wins and we need to have $\left.\kappa\right|_{s<0}=0$ as well as $\left.\kappa\right|_{s>0}=1$, proving part 2 (a) of the proposition.
@@
Next, assume voters generally believe that policy 0 will be implemented, independent of the true state $\omega$. Then $\kappa=0$ for all $s$. It follows that any voter with $s<0$ chooses $\tilde{\mu}^{*}$$\infty$, and thus all these voters hold belief $\pi(s,\mathbf{p},\tilde{\mu}^{*})=0$ and vote for policy 0. Moreover, any voter with $s>0$ chooses $\tilde{\mu}^{*}=0$, implying $\pi(s,\mathbf{p},\tilde{\mu}^{*})=\hat{\pi}$. Iff $\hat{\pi}\leq\tilde{\pi}$, this implies that indeed always a majority of voters votes for policy 0, and hence when $\hat{\pi}\leq\tilde{\pi}$ there exists an equilibrium of the voting subgame in which policy 0 is always chosen. This proves part 2 (b).
@@
Now assume voters generally believe that policy 1 will be implemented, independent of the true state $\omega$. Then $\kappa=1$ for all $s$, and any voter with $s>0$ chooses $\tilde{\mu}^{*}$$\infty$, and thus all these voters hold belief $\pi(s,\mathbf{p},\tilde{\mu}^{*})=1$ and vote for policy 1. Moreover, any voter with $s<0$ chooses $\tilde{\mu}^{*}=0$, inducing $\pi(s,\mathbf{p},\tilde{\mu}^{*})=\hat{\pi}$. If and only if $\hat{\pi}\geq\tilde{\pi}$, this implies that indeed always a majority of voters votes for policy 1, and hence when $\hat{\pi}\geq\tilde{\pi}$ there exists an equilibrium of the voting subgame in which policy 1 is always chosen. This proves part 2 (c).
@@
Finally, consider part 2 (d). An equilibrium in which the optimal policy never wins would imply policy 1 wins if $\omega=0$ and policy 0 wins if $\omega=1$. Hence, we would need to have $\kappa$ weakly decreasing in $s$. Consider a voter with signal $s>0$. It must be true that most of these voters vote for policy 0. Hence, it cannot be true that $\kappa\geq\tilde{\kappa}$. If $\kappa<\tilde{\kappa}$, these voters all hold belief $\hat{\pi}$. Next consider voters with $s<0$. It must be true that most of these voters vote for policy 1, implying $\kappa\leq\tilde{\kappa}$ is not possible. Thus, because $\kappa>\tilde{\kappa}$, these voters all also hold belief $\hat{\pi}$. If $\hat{\pi}>\tilde{\pi}$, policy 1 always wins, contradicting that the optimal policy never wins. If $\hat{\pi}<\tilde{\pi}$, policy 0 always wins, also contradicting that the optimal policy never wins. Finally, if $\hat{\pi}=\tilde{\pi}$, in each state, each policy wins with 50%, and thus also the optimal policy wins with a chance of 50%. This implies that the actual state is irrelevant for the probability of each policy winning, and thus we must have $\kappa=\frac{1}{2}$ for all $s$. But then it is not possible that $\kappa$ differs for voters with $s>0$ and voters with $s<0$. Hence, such an equilibrium of the voting subgame cannot exist.
This completes the proof of the proposition.∎
@@ A.5 Proof of Proposition 3
In a symmetric equilibrium, each candidate wins with a 50% probability. We next show that no lucrative deviations exist. Hence, suppose that off the equilibrium path voters observe $\tilde{\mathbf{p}}\in\{(0,1),(1,0)\}$.
$\Delta\geq 1$:
@@
We know from Corollary 1 that after $\pi(s,\tilde{\mathbf{p}},\tilde{\mu}^{*})=0$ if $s<0$ and $\pi(s,\tilde{\mathbf{p}},\tilde{\mu}^{*})=\hat{\pi}$ if $s\geq 0$. Hence, $\hat{\pi}\leq\tilde{\pi}$, and thus no voter ever votes for policy 1. It follows that policy 1 never wins. Consequently, a deviation from $\mathbf{p}=(0,0)$ is never profitable, and thus $\mathbf{p}^{*}=(0,0)$ is indeed an equilibrium. Moreover, $(1,1)$ cannot be an equilibrium, because a deviation is always profitable, because when spotting such a deviation, voters will ignore information implying that climate change is severe. Thus, the unique symmetric equilibrium is candidates choosing $\mathbf{p}^{*}=(0,0)$. This is supported by $\kappa^{*}=0$ and $\tilde{\mu}^{*}(s)$ as described in Lemma 1.
$\Delta<1$:
It is easy to show that the inefficient equilibrium with $\mathbf{p}^{*}=(0,0)$ still exists. Hence, here we focus on the existence of the efficient equilibrium.
Suppose candidates choose $\mathbf{p}=(1,1)$. Hence, as before, each wins with a 50% probability. If a candidate deviates, they face conflicting policy platforms, and hence do not learn from $\mathbf{p}$ itself. The only difference between voters is the signals they receive. Hence, if they hold different beliefs, these need to be based on signals.
[…]
@@
It follows from Proposition 1 part 2 (d) that no equilibrium can exist in which always the wrong policy is chosen. Hence, we now only need to show that no equilibrium, in which policy 1 always wins, can exist. Suppose to the contrary that voters expect policy 1 to be implemented with certainty, $\kappa=1\geq\tilde{\kappa}$ for all $s$. Then it follows from Corollary 1 that $\pi(s,\mathbf{p},\tilde{\mu})=q<\tilde{\pi}$ if $s<0$ and $\pi(s,\mathbf{p},\tilde{\mu})=1$ if $s>0$. Therefore, all voters vote informatively, implying $p=\omega$ wins if $p_{1}\neq p_{2}$, thus contradicting $\kappa=1$. Hence, it follows that no equilibrium, in which policy 1 is always chosen, exists.
This proves the proposition. ∎
Appendix B Costly Distortions
B.1 Cost from Imperfect Decision Making
Recall that voters aim at maximizing $W^{UC}=\lambda AU+(1-\lambda)C^{U}$. Each voter chooses $\tilde{\mu}$ to maximize $W$, taking into account that $\tilde{\mu}\neq\mu$ may negatively impact the quality of the decision at the ballot.
When $\tilde{\mu}$ is chosen in a way that is inconsequential for the decision at the ballot, $C^{U}=0$. Otherwise, $C^{U}$ equals the expected cost from deciding incorrectly.
If a voters chooses to vote for policy 1 at the ballot, her expected utility is
$u_{1}=\frac{-\Delta qe^{\frac{2\mu s}{\sigma^{2}}}+q-1}{q\left(e^{\frac{2\mu s}{\sigma^{2}}}-1\right)+1}.$
When voting for policy 0, expected utility is
$u_{0}=-\frac{q(b+\Delta)}{q-(q-1)e^{-\frac{2\mu s}{\sigma^{2}}}}$
Hence, a voter votes for policy 1 iff
$u_{1}\geq u_{0}\Leftrightarrow s>\tilde{s}:=\frac{\sigma^{2}\log\left(\frac{1-q}{bq}\right)}{2\mu}>0.$
Any voter with $s<\tilde{s}$ will vote for policy 0 with Bayesian information processing, and this does not change with motivated beliefs. Thus, If $s<\tilde{s}$, then $C^{U}=0$ because the choice at the ballot is identical to the Bayesian optimal choice. If $s>\tilde{s}$, this may change, as the Bayesian optimal choice would be to vote for policy 1, whereas with distorted information processing policy 0 might be chosen. If the choice at the ballot is indeed altered by choosing $\tilde{\mu}\neq\mu$, then the voter incurs a cost of
$C^{U}=-b+\frac{(b+1)(1-q)}{1-q+qe^{\frac{2\mu s}{\sigma^{2}}}}.$
This equals the expected utility loss from taking an incorrect decision.
As explained in the main text, when deciding if and how to distort beliefs, a voter has only two potentially optimal choices: $\tilde{\mu}=0$ and $\tilde{\mu}=\mu_{c}$, where $\mu_{c}$ is the distortion that induces a posterior belief of $1/(1+b)$. These lead to the following welfare levels: